OpenStudy (anonymous):

the vertical triangular plate shown below is the end plate of a feeding trough full of hog slop, weighing 80 pounds per cubic foot. what is the force against the plate?

OpenStudy (anonymous):



OpenStudy (anonymous):

is triangular plate conical in shape?

OpenStudy (anonymous):

it didn't specify whether or not the plate was conical. @surjithayer

OpenStudy (anonymous):

my point of asking was that if it is conical then i can find weight of hog slop replaced by cone, that is upthrust.

OpenStudy (anonymous):

volume of cone \[=\frac{ 1 }{ 3 } \pi 4^2*2=\frac{ 32 }{ 3 } \pi \] weight or upward force \[=\frac{ 32\pi }{3 }*80=?\]

OpenStudy (anonymous):

that answer is 214466. but the answer is supposed to be 426.67. how do i get that? @surjithayer

OpenStudy (anonymous):

would the shape be a triangular prism, perhaps?

OpenStudy (anonymous):

it is 2680.86 x is from -4 ft to 4ft?

OpenStudy (anonymous):

yes that's what the picture says

OpenStudy (anonymous):

i mean to say is it foot or something else.

OpenStudy (anonymous):

oh, feet sorry i forgot to put in the units

OpenStudy (anonymous):

i am sorry, now i have to go to bed. i will see tomorrow if new idea struck.